OpenStudy (anonymous):

find a quadratic equation whose roots are twice the roots of 2x^2-5x-3=0 ?

OpenStudy (anonymous):

first find the roots of the given euadratic equation

OpenStudy (anonymous):

its x=-1/2,x=3 then, ur required quadratic equation is (x+1/2)(x-3)

OpenStudy (anonymous):

(2x+1)(x-3)=0 ===> x=-1/2 or x=3 ===> the new roots are -1,6 ===> (x+1)(x-6)=0 ==> x^2+-5x-6=0